OpenStudy (anonymous):

Which of the following would be equivalent to 9^2 ⋅ 9^6?

OpenStudy (anonymous):

Which of the following would be equivalent to 92 ⋅ 96? 9 to the 8th power over 9 to the 7th power 92 ⋅ 97 9 to the 10th power over 9 to the 2nd power 90

OpenStudy (anonymous):

Do you remember your exponent rules? \[a^b\times a^c=a^{b+c}\]\[\frac{a^b}{a^c}=a^{b-c}\]Using these, try simplifying. So,\[9^2\times9^6=9^?\]